### v3

# New Factorizations of Yang-Mills Amplitudes

Abstract
We propose a new factorization pattern for tree-level Yang-Mills (YM) amplitudes, where they decompose into a sum of gluings of two lower-point amplitudes by setting specific two-point non-planar Mandelstam variables within a rectangular configuration to zero. This approach manifests the hidden zeros of YM amplitudes recently identified. Furthermore, by setting specific Lorentz products involving polarization vectors to zero, the amplitudes further reduce to a sum of products of three currents. These novel factorizations provide a fresh perspective on the structure of YM amplitudes, potentially enhancing our understanding and calculation of these fundamental quantities.
Introduction
Scattering amplitudes in quantum field theory factorize into sums of products of lower-point amplitudes at the residues of their poles, reflecting the fundamental property of locality. This factorization enables recursive computation of amplitudes, making it a powerful method in modern scattering amplitude research. A notable application of this concept is the Britto-Cachazo-Feng-Witten recursion Britto et al. (2005), which introduces an auxiliary variable in the complex plane to control amplitude factorization, leading to significant advancements.
At the loop level, (generalized) unitarity cuts are employed to determine the loop integrand of amplitudes Bern et al. (1994, 1995). The primary difference from conventional factorization is the substitution of a loop propagator for a tree propagator on-shell. Thus, studying factorizations also aids in understanding unitarity cuts at the loop level. Recently, factorizations have been utilized to determine Mellin amplitudes in AdS space (cf. Gonçalves et al. (2015); Alday et al. (2024); Cao et al. (2024a)).
Given the importance of factorizations in scattering amplitude studies, new types of factorizations have been investigated. In Cachazo et al. (2022), it was proposed that by setting specific Mandelstam variables to zero, amplitudes for biadjoint $\phi^{3}$ theory, the Nonlinear Sigma Model (NLSM), and a special Galileon theory Cachazo et al. (2015) split into three currents, a property termed semi-locality. This differs from conventional factorizations as no residues are computed  [^1] The smoothly splitting amplitudes for $\phi^{p}$ theories are discussed in Giménez Umbert and Yeats (2024). See more recent progress in Giménez Umbert and Sturmfels (2025); Early (2025).
In Arkani-Hamed et al. (2024a), another form of three-part splitting was discovered for $\text{Tr}(\phi^{3})$ theory and NLSM. By setting certain two-point Mandelstam variables in a rectangle to zero and then turning back on one of them, the amplitudes split into three parts near their zeros. These two types of three-part splitting for scalar theories are unified by a so-called 2-split proposed in Cao et al. (2024b, c), which asserts that amplitudes factorize into two currents under certain subspaces of the Mandelstam variables. Applying the 2-splits in different ways reproduces both types of three-part splitting found in Cachazo et al. (2022) and Arkani-Hamed et al. (2024a). Additionally, significant progress has been made in extending such split factorizations to loop orders by analyzing the curve integral formulation for ${\rm Tr}(\phi^{3})$ amplitudes, as discussed in Arkani-Hamed et al. (2024a, 2023a, b); Arkani-Hamed and Figueiredo (2024a, b).
Beyond scalar theories, hidden zeros in tree-level Yang-Mills (YM) amplitudes were also proposed in Arkani-Hamed et al. (2024a) by setting specific Mandelstam variables in a rectangle to zero along with their corresponding Lorentz products involving polarizations. Subsequently, the 2-split for YM and GR amplitudes was designed in Cao et al. (2024b, c) by turning on some vanishing Mandelstam variables while setting additional Lorentz products involving polarizations to zero. The hidden zeros in scalar theories have been further studied in Bartsch et al. (2024); Li et al. (2024) using double copy relations Kawai et al. (1986); Bern et al. (2008) and uniqueness bootstrap methods Rodina (2025), which are also claimed to be applicable to the hidden zeros in amplitudes of YM, GR, and other theories.
In this paper, we further investigate a new type of factorization of YM amplitudes by setting specific Mandelstam variables in a rectangle to zero, identical to the constraint used to produce zeros in certain scalar amplitudes Arkani-Hamed et al. (2024a). We found that under this constraint, YM amplitudes always factorize into sums of gluings of lower-point YM amplitudes. In addition to the summation over all possible states of the polarization of the internal gluon that is exchanged, as in conventional factorization, our factorization further includes a summation over different gluon pairs appearing in the vanishing Mandelstam variable rectangle.
[…]
Figure 1: New factorization of YM amplitudes: An $n$-point Yang-Mills amplitude factorizes into a sum over contributions, each involving a 3-point amplitude and an $(n-1)$-point amplitude. The sum runs over all permutations $\rho$ of the $(m-1)$ external gluons. The internal gluon $\hat{j}$ is exchanged in two diagrams. For simplicity, some rational functions of Mandelstam variables are omitted in this schematic representation.
The first graph in this figure also provides an intuitive explanation for the constraints on the Mandelstam variables in $h_{m}$, as defined in (1). Selecting any two non-adjacent external gluon legs (e.g., $m{+}1$ and $n$) in a color-ordered amplitude naturally partitions the remaining legs into two sets. Setting all two-particle Mandelstam variables that involve legs from both sets to zero gives $h_{m}=0$.
We refer to the contribution from each gluon pair as a gluon pair contribution. By eliminating all Lorentz products involving polarizations associated with the vanishing Mandelstam variables in the rectangle, we can observe how all gluon pair contributions vanish directly in our factorization formulas, thus making the hidden zeros manifest. Conversely, since the proposal of zeros in YM amplitudes in Arkani-Hamed et al. (2024a), an important question has been the behavior of amplitudes near these zeros when Lorentz products involving polarizations are turned on back. Our factorization formulas address this question precisely, demonstrating how contributions accumulate as more Lorentz products are restored, ultimately recovering the full factorization formulas.
Our factorization formulas for YM amplitudes differ from the 2-split proposed in Cao et al. (2024b, c) in that we only need to set certain Mandelstam variables to zero while allowing all Lorentz products involving polarization to remain. Additionally, our factorization formulas typically include a summation of gluing contributions instead of a single product of two amplitudes. However, combining our factorization formulas with the 2-split formulas yields another elegant result. By imposing conditions on Lorentz products such that only a single gluon pair contribution survives in our factorization summation for YM amplitudes, and then further eliminating additional Lorentz products involving polarizations, the entire YM amplitude factorizes into a sum of three currents. Continued study of our new factorization should uncover more properties of YM amplitudes, providing a novel description.
New Factorizations
Consider a matrix of Mandelstam variables $s_{ab}:=k_{a}\cdot k_{b}$, denoted as $h_{m}$ with $31\leq m\leq n-$ 3,
$h_{m}:=\begin{Bmatrix}s_{1\,m+2}\,,\,s_{1\,m+3}\,,\ldots,s_{1\,n-1}\,\,,\\ s_{2\,m+2}\,,\,s_{2\,m+3}\,,\ldots,s_{2\,n-1}\,\,,\\ \vdots\hskip 31.2982pt\vdots\hskip 48.36958pt\vdots\\ \hskip 8.5359pts_{m\,m+2},s_{m\,m+3},\ldots,s_{m\,n-1}~{}~{}\end{Bmatrix}\,.$ (1)
As established in Arkani-Hamed et al. (2024a), setting $h_{m}=0$  [^2] We assume the spacetime dimension is general enough to ignore the Gram determinant constraints. results in the vanishing of $n$-point doubly color-ordered biadjoint $\phi^{3}$ amplitudes with two canonical orderings, as well as many other scalar amplitudes. YM amplitudes do not vanish under this constraint; in this paper, we analyze how they factorize.
Main Claim
We claim that the $n$-point YM amplitudes with canonical ordering factorize into sums of gluings of a three-point and an $(n-1)$-point YM amplitude in the subspace of Mandelstam variables $h_{m}=0$ for $m\geq 1$ and $n\geq m+3$,
$A({\mathbb{I}}_{n})\!\xrightarrow{h_{m}=0}\!\!\!\!\sum_{\begin{subarray}{c}1\leq i\leq m<m+2\leq j\leq n-1\end{subarray}}\!\!\!F_{m,n}(i,j)\,,$ (2)
$F_{m,n}(i,j)=\frac{(-1)^{n+1}}{s_{12\ldots m+1}}\sum_{\epsilon_{\hat{j}}}A(ij-\!\!\hat{j})\sum_{\rho\in S_{m-1}}X(s,\rho)$
×$\quad\times A\big{(}\rho(12\ldots i\!-\!1\,i\!+\!1\ldots m)\,m{+}1\,m{+}2\ldots{\hat{j}}\ldots n\big{)}\,.$ (3)
This factorization, illustrated in fig. 1, involves summing over all gluon pairs $(i,j)$ appearing in the matrix $h_{m}$. Each term in the sum, denoted $F_{m,n}(i,j)$, represents a “gluon pair contribution”. Additionally, we sum over all possible polarization states of the internal gluon $\hat{j}$, which is exchanged between the two amplitudes.
Momentum conservation in the three-point amplitude $A(ij-\hat{j})$ ensures that the internal gluon satisfies $k_{\hat{j}}=-k_{-\hat{j}}=k_{i}{+}k_{j}$. Under the constraint $h_{m}=0$, the Mandelstam variable $s_{ij}$ vanishes, making the exchanged gluon on-shell. Finally, we sum over all permutations $\rho$ of the $m{-}1$ gluons $\{1,2,\dots,i{-}1,i{+}1,\dots,m\}$.
The terms $X(s,\rho)$ are ratios of Mandelstam variables that involve only the gluons $11,2,\ldots,m+$ 1 and they remain independent of the total number of particles $n$. Examples of $X(s,\rho)$ for low values of $m$ are as follows,
$\text{For }m=1,i=1,\quad X(s,\emptyset)=1\,,$
$\text{For }m=2,i=1,\quad X(s,2)=\frac{s_{12}+s_{23}}{s_{12}}\,,$
$\text{For }m=2,i=2,\quad X(s,1)=-\frac{s_{13}}{s_{12}}\,.$ (4)
In general, $X(s,\rho)$ is expressed as
[…]
×$\times\,{\cal B}_{m+1,i}[12\ldots i{-}1,i{+1}\ldots m|\pi]\,,$ (5)
where $g^{-1}$ is the inverse of the matrix $g[\rho,\pi]$, defined by
$g[\rho,\pi]:=$ $\sum_{\begin{subarray}{c}a=1\\ a\neq i\end{subarray}}^{m+1}\left(\sum_{b\in Y(a,\rho)}s_{ib}+s_{i\,m+1}\right)$
×$\times{\cal B}_{m+1,i}[Y(\rho,a),Y(a,\rho)|\pi]\,.$ (6)
[…]
The term ${\cal B}_{m+1,i}[\alpha,\beta|\pi]$ in (Main Claim) and (Main Claim) is derived from the coefficients in the BCJ relations for YM amplitudes, as outlined in Bern et al. (2024, 2008),
[…]
Here, $k_{\hat{n}}$, determined by momentum conservation, is typically on-shell to maintain gauge invariance in YM amplitudes. However, in this context, the on-shell condition for $k_{\hat{n}}$ is relaxed while ensuring that ${\cal B}_{m+1,i}[\alpha,\beta|\pi]$ remains independent of $k_{\hat{n}}$.
It turns out that the denominators of $X(s,\rho)$ correspond to all planar poles involving gluon $i$, ranging from $1$ to $m$. Numerically, this structure has been verified up to $m=7$. The numerators are polynomials in the Mandelstam variables $s_{rt}$ for $11\leq r<t\leq m+$ 1 and depend on the ordering $\rho$. Explicit forms of ${X}(s,\rho)$ with polynomial numerators for $m\leq 5$ are provided in ancillary files.
Finally, the contributions $F_{m,n}(i,j_{1})$ and $F_{m,n}(m+1-i,j_{2})$, where $j_{1},j_{2}=m+2,m+3,\ldots,n-1$ and $j_{1}+j_{2}=n+m+1$, are equivalent up to a sign under the relabeling $a\leftrightarrow m+1-a$. As a result, $X(s,\rho)$ satisfies the symmetry,
$X(s,\rho^{T}\big{|}_{a\to m+1-a\,{\rm mod}\,n})=X(s,\rho)\big{|}_{{a\to m+1-a\,{\rm mod}\,n}}\,,$ (8)
on the support of $h_{m}=0$, i.e. $s_{rn}=-\sum_{t=1,t\neq r}^{m+1}s_{rt}$ for $1\leq r\leq m$.
As an illustration, in (4), the term $X(s,1)$ can be rewritten as $\frac{s_{12}+s_{1n}}{s_{12}}$ on the support of $h_{2}=0$, making the symmetry (8) explicit when compared with the expression for $X(s,2)$.
A rigorous proof of the new factorization formula (2) is provided in the companion paper Zhang (2025).
Properties
Applying the completeness relation of the polarization vector to our formula (3), we get a product of currents
$\sum_{\epsilon_{\hat{j}}}\!A(ij-\!\!\hat{j})A\big{(}\rho\ldots{\hat{j}}\ldots n\big{)}\!=\!A_{\mu}(ij;-\hat{j})A^{\mu}(\ldots n\,\rho\ldots;\hat{j}),$ (9)
with
$A^{\mu}(ij;-\hat{j})=\epsilon_{i}\!\cdot\!\epsilon_{j}\,k_{i}^{\mu}+\epsilon_{i}\!\cdot\!k_{j}\,\epsilon_{j}^{\mu}-k_{i}\!\cdot\!\epsilon_{j}\,\epsilon_{i}^{\mu}\,.$ (10)
Define $x_{ij}:=\{\epsilon_{i}\cdot\epsilon_{j},\epsilon_{i}\cdot k_{j},k_{i}\cdot\epsilon_{j}\}$ and denote their collection as
$\hat{H}_{m}:=\begin{Bmatrix}x_{1\,m+2}\,,\,x_{1\,m+3}\,,\ldots,x_{1\,n-1}\,\,,\\ x_{2\,m+2}\,,\,x_{2\,m+3}\,,\ldots,x_{2\,n-1}\,\,,\\ \vdots\hskip 31.2982pt\vdots\hskip 48.36958pt\vdots\\ \hskip 8.5359ptx_{m\,m+2},x_{m\,m+3},\ldots,x_{m\,n-1}~{}~{}\end{Bmatrix}\,.$ (11)
Additionally, denote $H_{m}:=h_{m}\sqcup\hat{H}_{m}$. Setting $x_{ij}=0$ obviously results in $A^{\mu}(ij;-\hat{j})=0$ as shown in (10), hence $F_{m,n}(i,j)=0$ following (3). When setting all $x_{ij}\in\hat{H}_{m}$ to zero, the entire amplitude $A(\mathbb{I}_{n})$ vanishes according to (2),
$A({\mathbb{I}}_{n})\!\xrightarrow{H_{m}=0}0\,,$ (12)
recovering the hidden zeros of the YM amplitudes as proposed in Arkani-Hamed et al. (2024a). Our factorization formulas make this property manifest.
Step backward. When restoring a single $x_{ij}\neq 0$ from the constraint $H_{m}=0$, the whole amplitude simplifies to a single gluon pair contribution,
$A({\mathbb{I}}_{n})\!\xrightarrow{(H_{m}/\{x_{ij}\})=0}F_{m,n}(i,j)\,.$ (13)
This provides a physical meaning for the function $F_{m,n}(i,j)$. Compared with the conventional factorization, our new factorization (2) has an additional summation over the gluon pairs $(i,j)$. However, for each pair, we can assume that all other $x_{i^{\prime}j^{\prime}}\neq x_{ij}\in\hat{H}_{m}$ vanish and consider its contribution individually. Then, their direct addition provides the whole factorization under the constraint $h_{m}=0$.
When only one element, say $\epsilon_{i}\cdot k_{j}\in x_{ij}$, is turned on instead of the whole $x_{ij}$, only one term in the current $A_{\mu}(ij;-\hat{j})$ given in (10) survives, and the YM amplitude reduces to a combination of $(n-1)$-point amplitudes,
$A(\mathbb{I}_{n})$ ${\rm except}~{}\epsilon_{i}\cdot k_{j}\neq 0]{H_{m}=0}\frac{\epsilon_{i}\cdot k_{j}}{s_{12\ldots m+1}}\sum_{\rho\in S_{m-1}}X(s,\rho)$ (14)
×$\times A\big{(}\rho(12\ldots i\!-\!1\,i\!+\!1\ldots m)\,m{+}1\,m{+}2\ldots{\hat{j}}\ldots n\big{)}\,,$
where the momentum $k_{\hat{j}}=k_{i}+k_{j}$ and polarization $\epsilon_{\hat{j}}=\epsilon_{j}$ of the new gluon $\hat{j}$ are on-shell and orthogonal to each other on the support of the conditions $H_{m}=0$ except $\epsilon_{i}\cdot k_{j}\neq 0$. If only $k_{i}\cdot\epsilon_{j}\in x_{ij}$ is turned on, $\epsilon_{\hat{j}}=-\epsilon_{i}$. When turning on $\epsilon_{i}\cdot\epsilon_{j}$ instead, $\epsilon_{\hat{j}}$ can take either $k_{i}$ or $-k_{j}$, reflecting the gauge redundancy of the amplitudes.
Following the discovery of zeros of YM amplitudes as described in Arkani-Hamed et al. (2024a), a prompt inquiry has been the response of amplitudes near these zeros when the Lorentz products involving polarizations are turned back on. Our factorization formulas provide a precise answer to this question.
In the following, we show more details for the cases $m=1,2,3,4,5$ respectively.
Cases $m=1,2$
The simplest application of the new factorization (2) is for $m=1$,
$A({\mathbb{I}}_{n})\xrightarrow{s_{13}=s_{14}\ldots=s_{1\,n\!-\!1}=0}$ (15)
$\quad\frac{(-1)^{n+1}}{s_{12}}\sum_{j=3}^{n-1}\sum_{\epsilon_{\hat{j}}}A(1j\,-\!\hat{j})A(23\ldots j\!-\!1\,\hat{j}\,j\!+\!1\ldots n)\,.$
Two quick examples are given by
$A(1234)$ $\xrightarrow{s_{13}=0}-\frac{1}{s_{12}}\sum_{\epsilon_{\hat{3}}}A(13-\!\hat{3})A(2\hat{3}4)\,,$
$A(12345)$ $\xrightarrow{s_{13}=s_{14}=0}\frac{1}{s_{12}}\big{(}\sum_{\epsilon_{\hat{3}}}A(13-\!\hat{3})A(2\hat{3}45)$
$\qquad\qquad+\sum_{\epsilon_{\hat{4}}}A(14-\!\hat{4})A(23\hat{4}5)\big{)}\,.$ (16)
For $m=2$ and $n\geq 5$, we have,
$A({\mathbb{I}}_{n})\xrightarrow[s_{24}=s_{25}\ldots=s_{2\,n\!-\!1}=0]{s_{14}=s_{15}\ldots=s_{1\,n\!-\!1}=0}$ (17)
$\quad(-1)^{n+1}\frac{s_{12}+s_{23}}{s_{12}s_{123}}\sum_{j=4}^{n-1}\sum_{\epsilon_{\hat{j}}}A(1j\,-\!\hat{j})A(234\ldots\hat{j}\ldots n)$
$\quad+(-1)^{n}\frac{s_{13}}{s_{12}s_{123}}\sum_{j=4}^{n-1}\sum_{\epsilon_{\hat{j}}}A(2j\,-\!\hat{j})A(134\ldots\hat{j}\ldots n)\,,$
where $k_{\hat{j}}=k_{1}+k_{j}$ and $k_{2}+k_{j}$ respectively in the last two lines implied by the momentum conservation of the three-point amplitudes in their lines.
Here we provide an example for $n=5$,
$A(12345)$ $\xrightarrow{s_{14}=s_{24}=0}\frac{s_{12}+s_{23}}{s_{12}s_{123}}\sum_{\epsilon_{\hat{4}}}A(14-\!\hat{4})A(23\hat{4}5)$
$\quad-\frac{s_{13}}{s_{12}s_{123}}\sum_{\epsilon_{\hat{4}}}A(24-\!\hat{4})A(13\hat{4}5).$ (18)
Obviously, $A(12345)\big{|}_{s_{13}=s_{14}=0}$ and $A(12345)\big{|}_{s_{14}=s_{24}=0}$ are identified by relabeling $a\leftrightarrow a^{\prime}=5-a$ modulo $5$ for all $1\leq a\leq 5$, leading to non-trivial relations in their reductions on the right-hand sides of (Cases $m=1,2$) and (Cases $m=1,2$), but they match after relabeling under the given conditions $s_{13}=s_{14}=0$.
Cases $m=3,4,5$
For $m=3$, the gluon pair contributions $F_{3,n}(i,j)$ with $i=1,2,3$ are given by
$F_{3,n}(1,j)$ $=\frac{(-1)^{n+1}}{s_{12}s_{123}s_{1234}}\sum_{\epsilon_{\hat{j}}}A(1j-\!\hat{j})$
[…]
$-A(3245\ldots{\hat{j}}\ldots n)s_{3n}s_{24}\Big{)}\,,$ (19)
$F_{3,n}(2,j)$ $=\frac{(-1)^{n+1}}{s_{12}s_{23}s_{123}s_{1234}}\sum_{\epsilon_{\hat{j}}}A(2j-\!\hat{j})$
[…]
$\qquad\qquad+s_{12}s_{23}(s_{13}+s_{14}+s_{3n})\big{)}$
$+A(3145\ldots{\hat{j}}\ldots n)s_{14}s_{3n}(s_{12}+s_{23})\,\Big{)}\,,$ (20)
$F_{3,n}(3,j)$ $=\frac{(-1)^{n+1}}{s_{23}s_{123}s_{1234}}\sum_{\epsilon_{\hat{j}}}A(3j-\!\hat{j})$
[…]
$-A(2145\ldots{\hat{j}}\ldots n)s_{14}s_{2n}\Big{)}\,.$ (21)
Summing over all such contributions gives the factorization of YM amplitudes under the condition $h_{3}=0$, as shown in (2). The explicit forms of $X(s,\rho)$ for $m=3$ can be directly derived from these contributions. For example, from the final term in (20), we find $X(s,31)=\frac{s_{14}s_{3n}(s_{12}+s_{23})}{s_{12}s_{23}s_{123}}$ for $F_{3,n}(2,j)$. Furthermore, it can be verified that $X(s,13)$ satisfies the symmetry property in (8) under the relabeling $a\leftrightarrow a^{\prime}=n+4-a\,{\rm mod}\,n$ for all $1\leq a\leq n$.
For $m=4$, each gluon pair contribution $F_{4,n}(i,j)$ involves six $(n{-}1)$-point amplitudes, $A(\rho(12\ldots i\!\!/\ldots 4)56\ldots\hat{j}\ldots n)$ with $\rho\in S_{3}$. Their explicit expressions are present in the supplemental materials. All ${X}(s,\rho)$ for $m\leq 5$ are also provided in ancillary files.
Further Factorizations
Turning on a single $x_{ij}$ from the constraint $H_{m}=0$, the factorization of the YM amplitudes is described by a single gluon pair contribution $F_{m,n}(i,j)$ given in (3). For $m\geq 2$, there are still many vanishing Lorentz products in the $(n-1)$ point amplitudes $A\big{(}\rho(12\ldots i\!-\!1\,i\!+\!1\ldots m)\,m{+}1\,m{+}2\ldots{\hat{j}}\ldots n\big{)}$,
$s_{ab}$ $=\epsilon_{a}\cdot k_{b}=k_{a}\cdot\epsilon_{b}=\epsilon_{a}\cdot\epsilon_{b}=0,$
$\forall$ $a\in\{1,2,\ldots,i\!-\!1,i\!+\!1,\ldots,m\},$
$b\in\{m{+}2,m\!+\!3,\ldots j\!-\!1,j\!+\!1,\ldots,n\!-\!1\}\,.$
Thus, we consider further possible reductions of the amplitudes. For Lorentz products involving $k_{\hat{j}}=k_{i}+k_{j}$, we have
$s_{a\hat{j}}=k_{a}\cdot k_{\hat{j}}=k_{a}\cdot k_{i}\neq 0,\quad\epsilon_{a}\cdot k_{\hat{j}}=\epsilon_{a}\cdot k_{i}\neq 0\,.$
Although $s_{aj}\in h_{m}=0$, $s_{a\hat{j}}$ does not vanish, prompting the further factorization of the $(n-1)$-point amplitudes $A\big{(}\rho\ldots\hat{j}\ldots n\big{)}$ into a product of two currents based on the 2-split construction proposed in Cao et al. (2024b, c), with additional Lorentz products involving polarizations needing to be set to zero.
These Lorentz products include $\epsilon_{a}\cdot\epsilon_{b^{\prime}}$ and $k_{a}\cdot\epsilon_{b^{\prime}}$ with $b^{\prime}\in\{\hat{j},m+1,n\}$. However, after summing over all states of $\epsilon_{\hat{j}}$ according to (9), it is $\epsilon_{j}^{\mu},k_{j}^{\mu}$, or $\epsilon_{i}^{\mu}$ in $A^{\mu}(ij;-\hat{j})$ that contract with the current $A^{\mu}(\rho\ldots{\hat{j}}\ldots n)$. Hence, $\epsilon_{j}^{\mu},k_{j}^{\mu}$, or $\epsilon_{i}^{\mu}$ play the role of $\epsilon_{\hat{j}}$ after the state sum. Correspondingly, we examine their contractions with $\epsilon_{a}$ or $k_{a}$ and get
$k_{a}\cdot\epsilon_{j}=k_{a}\cdot k_{j}=0,\quad k_{a}\cdot\epsilon_{i}\neq 0\,,$
$\epsilon_{a}\cdot\epsilon_{j}=\epsilon_{a}\cdot k_{j}=0,\quad\epsilon_{a}\cdot\epsilon_{i}\neq 0\,.$
Therefore, based on the 2-split construction proposed in Cao et al. (2024b, c), we just need to further impose the conditions
$\epsilon_{a}\cdot\epsilon_{b^{\prime}}=k_{a}\cdot\epsilon_{b^{\prime}}=0\,,$ (22)
$\quad\forall a\in\{1,2,\ldots,i\!-\!1,i\!+\!1,\ldots,m\},\quad b^{\prime}\in\{i,m{+}1,n\}\,,$
so that all contractions between $\epsilon_{j}^{\mu},k_{j}^{\mu}$, or $\epsilon_{i}^{\mu}$ and $A^{\mu}(\rho\ldots{\hat{j}}\ldots n)$ split into two currents.
In this way, we achieve a three-split of the YM amplitude for $32\leq m\leq n-$ 3, where it decomposes into a sum of products of three currents,
$A(\mathbb{I}_{n})$ (22)${(H_{m}/\{x_{ij}\})=0}\frac{(-1)^{n+1}}{s_{12\ldots m+1}}A^{\mu}(ij;-\hat{j})\Big{[}\sum_{\rho\in S_{m-1}}X(s,\rho)$ )
×$\times J^{{\rm YM}+\phi^{3}}(n^{\phi}\rho(12\ldots i\!-\!1\,i\!+\!1\ldots m)\,(m{+}1)^{\phi}\,;\tilde{j}^{\prime}{}^{\phi})\Big{]}$
×$\times J^{\rm YM}_{\mu}(j\!+\!1\,j\!+\!2\ldots n\,m\!+\!1\,m\!+\!2\,\ldots j\!-\!1;\tilde{j})\,.$ (23)
Here, $J^{{\rm YM}+\phi^{3}}(n^{\phi}\ldots;\tilde{j}^{\prime}{}^{\phi})$ are single-trace currents for a mixed theory, dubbed as ${\rm YM}+\phi^{3}$, with two on-shell scalars $n,m+1$ and one off-shell scalar $\tilde{j}^{\prime}{}^{\phi}$. $J^{\rm YM}(j+1\ldots j-1;\tilde{j})$ is a YM current with an off-shell leg $\tilde{j}$. The momenta of the off-shell particles $\tilde{j}^{\prime}{}^{\phi}$ and $\tilde{j}$ are constrained by momentum conservation in their currents. In practice, one can use lower-point amplitude formulas to compute these currents, ensuring that all momenta of the off-shell particles including those in the Mandelstam variables and $s_{m\!+\!1\,n}$ are eliminated on the support of momentum conservation in each current. For more details on the currents, refer to Cao et al. (2024b, c); Naculich (2015).
Here is an example of (Further Factorizations),
$A(12345)$ ${\begin{subarray}{c}\epsilon_{2}\cdot\epsilon_{1}=\epsilon_{2}\cdot\epsilon_{3}=\epsilon_{2}\cdot\epsilon_{5}=0\\ k_{2}\cdot\epsilon_{1}=k_{2}\cdot\epsilon_{3}=k_{2}\cdot\epsilon_{5}=0\end{subarray}}]{s_{14}=s_{24}=0,x_{24}=0}\frac{s_{12}+s_{23}}{s_{12}s_{123}}$ (24)
×$\times A^{\mu}(14;-\hat{4})J^{{\rm YM}+\phi^{3}}(5^{\phi}23^{\phi};\tilde{4}^{\prime}{}^{\phi})J^{\rm YM}_{\mu}(53;\tilde{4})\,,$
where $J^{{\rm YM}+\phi^{3}}(5^{\phi}23^{\phi};\tilde{4}^{\prime}{}^{\phi})=-\epsilon_{2}\cdot k_{3}/s_{23}+\epsilon_{2}\cdot k_{5}/s_{25}$ and $J^{\rm YM}_{\mu}(53;\tilde{4})=\epsilon_{5}\!\cdot\!\epsilon_{3}\,k_{5}^{\mu}+\epsilon_{5}\!\cdot\!k_{3}\,\epsilon_{3}^{\mu}-k_{5}\!\cdot\!\epsilon_{3}\,\epsilon_{5}^{\mu}$.
Discussion
In this paper, we investigated a novel factorization of YM amplitudes by setting specific non-planar two-point Mandelstam variables in a rectangular configuration to zero, which is the same constraint used to reveal zeros in some scalar theories as proposed in Arkani-Hamed et al. (2024a). Our new factorization formulas involve summing over gluon pairs appearing in the vanishing Mandelstam variable matrix. By setting all Lorentz products involving polarizations associated with the vanishing Mandelstam variables to zero, we observed how the contribution from each gluon pair in our factorization summation cancels out one by one, ultimately revealing the hidden zeros of YM amplitudes. In cases where only the contribution from a single gluon pair survives, setting additional Lorentz products involving polarizations to zero reduces the amplitudes into a sum of products of three currents. This work introduces a novel type of factorization, distinct from conventional methods that rely on residues at poles, and opens the possibility of recursively constructing amplitudes using our new factorization formulas.
In the companion paper Zhang (2025), we will provide proof of our new factorization formulas using the Cachazo-He-Yuan formalism Cachazo et al. (2014a). Under the constraint of vanishing Mandelstam variables, one can prove that all solutions to the scattering equations Cachazo et al. (2014b, c) become singular, making it evident why the amplitudes for YM theory (and many other theories such as GR, in addition to scalar theories) vanish when certain Lorentz products are set to zero. We will then demonstrate how these singular solutions independently contribute to yield new factorization formulas for YM theories when Lorentz products involving polarizations are turned back on.
Returning to our current work, it would be highly valuable to interpret the ratios of Mandelstam variables $X(s,\rho)$ in (Further Factorizations) as amplitudes or currents associated with specific theories, or even to uncover a Feynman rule that generates them. Exploring the new factorization at the level of Feynman diagrams offers another promising direction. This perspective could shed light on how each summand in the factorization operates independently within different Feynman diagrams, providing deeper insights into the physical significance of the hidden zeros in YM amplitudes Rodina (2025); Zhou (2024); Huang et al. (2025)  [^4] See recent progress at loop level in Backus and Rodina (2025).
While this paper focuses on YM amplitudes, it is equally valuable to investigate potential new factorizations for other theories, such as GR, and their extensions in string theories, both bosonic and supersymmetric Green et al. (1988); Mafra et al. (2013); He and Zhang (2017); He et al. (2019). By examining the behavior of their amplitudes around the zeros, we could derive new factorization formulas for these theories. Additionally, exploring their connections to double copy constructions Kawai et al. (1986); Bern et al. (2008) and soft theorems Weinberg (1965) would be highly beneficial. Extending this study from tree-level to loop-level amplitudes is another promising direction. By identifying zeros in loop integrands, we can further explore their new factorization properties. Further insights can be gained from studying the curve integral formalism for ${\rm Tr}(\phi^{3})$ amplitudes, which provides a framework for deriving YM amplitudes at all loop levels via the so-called “scalar scaffolding” procedure Arkani-Hamed et al. (2024a, 2023b, 2023c, 2023a, b); Arkani-Hamed and Figueiredo (2024a, b); De et al. (2024); Arkani-Hamed et al. (2024c).
Acknowledgements.
We are deeply grateful to Freddy Cachazo for suggesting this study, providing invaluable guidance during its early stages, and offering continued support and encouragement throughout the project. Our thanks also go to Qu Cao, Jin Dong, Song He and Bruno Umbert for their insightful discussions and constructive feedback on our manuscript. The research of Y.Z. was supported in part by a grant from the Gluskin Sheff/Onex Freeman Dyson Chair in Theoretical Physics and by Perimeter Institute. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities.
Supplemental materials
Appendix A New factorization for $m=4$
For $m=4$, the gluon pair contributions $F_{4,n}(i,j)$ with $i=1,2$ are given by
$F_{4,n}(1,\,$ $j)=\frac{1}{s_{12}s_{123}s_{1234}s_{12345}}\sum_{\epsilon_{\hat{j}}}A(1j-\!\hat{j})$ (25)
[…]
$+$ $A(24356\ldots{\hat{j}}\ldots n)s_{35}s_{2n}s_{4n}$
$+$ $A(32456\ldots{\hat{j}}\ldots n)(s_{24}\!+\!s_{25})s_{3n}(s_{4n}\!-\!s_{123})$
$+$ $A(34256\ldots{\hat{j}}\ldots n)s_{25}s_{3n}(s_{4n}\!-\!s_{123})$
-- $A(42356\ldots{\hat{j}}\ldots n)(s_{12}\!+\!s_{23}\!+\!s_{25})s_{35}s_{4n}$
$+$ $A(43256\ldots{\hat{j}}\ldots n)s_{25}(s_{12}\!-\!s_{35})s_{4n}\Big{)}\,,$
$F_{4,n}(2,\,$ $j)=\frac{1}{s_{12}s_{23}s_{123}s_{234}s_{1234}s_{12345}}\sum_{\epsilon_{\hat{j}}}A(2j-\!\hat{j})$ (26)
[…]
$\qquad\qquad\!-\!s_{12}^{2}((s_{13}s_{23}\!+\!s_{14}(s_{23}\!+\!s_{34}\!+\!s_{35}))s_{234}\!+\!s_{15}(s_{23}\!+\!s_{34}\!+\!s_{35})(s_{45}\!+\!s_{234}))$
$\qquad\qquad\!-\!s_{12}\big{(}2s_{23}s_{234}s_{13}^{2}\!+\!s_{13}(s_{15}((s_{34}\!+\!s_{35})(s_{45}\!+\!s_{234})\!+\!s_{23}(s_{45}\!+\!2s_{234}))$
$\qquad\qquad\!+\!s_{234}(s_{14}(3s_{23}\!+\!s_{34}\!+\!s_{35})\!+\!s_{23}(s_{23}\!+\!s_{234}\!+\!s_{345})))$
$\qquad\qquad\!+\!(s_{23}\!+\!s_{34}\!+\!s_{35})((s_{14}\!+\!s_{15})s_{234}^{2}\!+\!(s_{14}\!+\!s_{15})(s_{14}\!+\!s_{23}\!+\!s_{45})s_{234}\!+\!s_{15}s_{23}s_{45})\big{)}\big{)}$
-- $A(14356\ldots{\hat{j}}\ldots n)s_{35}(s_{12}(s_{14}s_{234}(\!-\!s_{4n}\!+\!s_{13}\!+\!s_{23})\!+\!s_{15}(s_{14}s_{234}\!+\!(s_{24}\!+\!s_{34}\!+\!s_{45})(s_{123}\!+\!s_{234})))$
$\qquad\qquad\!+\!s_{23}s_{234}((s_{13}\!+\!s_{15})(s_{14}\!+\!s_{24}\!+\!s_{34}\!+\!s_{45})\!-\!s_{14}s_{4n})\!+\!s_{14}s_{234}s_{12}^{2})$
$+$ $A(31456\ldots{\hat{j}}\ldots n)s_{3n}((s_{14}\!+\!s_{15})s_{234}(s_{23}(\!-\!s_{4n}\!+\!s_{13}\!+\!s_{23})\!+\!s_{12}^{2}\!+\!(s_{13}\!+\!s_{14}\!+\!s_{23}\!+\!s_{45})s_{12})$
$\qquad\qquad\!+\!s_{12}(s_{14}\!+\!s_{15})s_{234}^{2}\!+\!s_{12}s_{15}s_{45}s_{123})$
$+$ $A(34156\ldots{\hat{j}}\ldots n)s_{15}s_{3n}(s_{23}s_{234}(\!-\!s_{4n}\!+\!s_{13}\!+\!s_{23})\!+\!s_{12}(s_{123}\!+\!s_{234})(s_{23}\!-\!s_{4n}))$
$+$ $A(41356\ldots{\hat{j}}\ldots n)s_{35}s_{4n}((s_{13}\!+\!s_{15})s_{23}s_{234}\!+\!s_{12}s_{15}(s_{123}\!+\!s_{234}))$
$+$ $A(43156\ldots{\hat{j}}\ldots n)s_{15}s_{4n}(s_{23}s_{35}s_{234}\!+\!s_{12}(s_{13}\!+\!s_{35})(s_{123}\!+\!s_{234}))\Big{)}\,.$
Performing a relabeling $a\to a^{\prime}=n+5-a$ modulo $n$ for all $1\leq a\leq n$ on (25) and (26), we obtain $F_{4,n}(4,j)$ and $F_{4,n}(3,j)$, respectively. Summing over such functions provides the factorization of YM amplitudes under $h_{4}=0$ according to (2).